# Groupoid Graded Semisimple Rings

### v1

@@ 3.6. Some isomorphisms between categories of graded modules
is an isomorphism of groupoids, see for example [6, p.125]. Of course, if $\Gamma$ is a groupoid of the form $I\times G\times I$ for some group $G$ and set $I$, there is a natural choice of isomorphism in (3.1) making $\sigma_{e}=(e_{0},1_{G},e)$ for all $e\in I\equiv\Gamma_{0}$. We remark on passing that we obtain a relation between rings graded by a connected groupoid and categories graded by a group from Example 2.3(2) and (3.1).
Keep in mind the context of the previous paragraph and let $R=\bigoplus_{\gamma\in\Gamma}R_{\gamma}$ be a $\Gamma$-graded ring. We proceed to define some categories of right $R$-modules.
The objects of the category $G-\operatorname{gr\textrm{-}\!}R$ are the $G$-graded $R$-modules. That is, right $R$-modules $M$ such that $MR=M$ and for which there exists a family $\{M_{g}\colon g\in G\}$ of additive subgroups of $M$ such that $M=\bigoplus_{g\in G}M_{g}$ as additive groups and $M_{g}R_{\gamma}\subseteq M_{gg_{\gamma}}$ for each $\gamma\in\Gamma$. If $N$ is another $G$-graded right $R$-module, a homomorphism of $G$-graded modules is a homomorphism of modules $p\colon M\rightarrow N$ such that $p(M_{g})\subseteq N_{g}$ for all $g\in G$. An example of $G$-graded $R$-module is $R(e)$, where $e\in\Gamma_{0}$, via $R(e)_{g}:=\bigoplus_{f\in\Gamma_{0}}R_{\sigma_{e}^{-1}g\sigma_{f}}$ for each $g\in G$.
The objects of the category $(G\times\Gamma_{0})-\operatorname{gr\textrm{-}\!}R$ are the $(G\times\Gamma_{0})$-graded $R$-modules $M$. That is, right $R$-modules $M$ such that $MR=M$ and for which there exists a family of additive subgrups $\{M_{(g,e)}\colon(g,e)\in G\times\Gamma_{0}\}$ such that $M=\bigoplus\limits_{(g,e)\in G\times\Gamma_{0}}M_{(g,e)}$ as additive groups and, for each $\gamma\in\Gamma$, $M_{(g,e)}R_{\gamma}\subseteq M_{(gg_{\gamma},d(\gamma))}$ if $r(\gamma)=e$ and $M_{(g,e)}R_{\gamma}=0$ otherwise. Given $(G\times\Gamma_{0})$-graded $R$-modules $M,N$, a homomorphism of $(G\times\Gamma_{0})$-graded modules is a homomorphism of $R$-modules $p\colon M\rightarrow N$ such that $p(M_{(g,e)})\subseteq N_{(g,e)}$ for all $(g,e)\in G\times\Gamma_{0}$. An example of $(G\times\Gamma_{0})$-graded $R$-module is $R(e)$, where $e\in\Gamma_{0}$, via $R(e)_{(g,f)}:=R_{\sigma_{e}^{-1}g\sigma_{f}}$ for each $(g,f)\in G\times\Gamma_{0}$.
@@
A $G$-graded right $R$-module $M=\bigoplus_{g\in G}M_{g}$ has a natural structure of $(G\times\Gamma_{0})$-graded module. Indeed, for each $g\in G$ and $f\in\Gamma_{0}$, define
@@ 3.7. Gradings on retangular matrix groups
Let $I,J$ be nonempty sets. We denote by $\operatorname{M}_{I\times J}(R)$ the additive group consisting of the $I\times J$ matrices with entries in $R$ and with at most a finite number of nonzero entries. Suppose, moreover, that $\operatorname{supp}(R)\subseteq e\Gamma e$ for some $e\in\Gamma_{0}$. If $\overline{\sigma}=(\sigma_{i})_{i\in I}\in(e\Gamma)^{I},$ $\overline{\tau}=(\tau_{j})_{j\in J}\in(e\Gamma)^{J}$. We will denote by $\operatorname{M}_{I\times J}(R)(\overline{\sigma})(\overline{\tau})$ the $\Gamma$-graded additive group whose homogeneous component of degree $\gamma\in\Gamma$, $\operatorname{M}_{I\times J}(R)(\overline{\sigma})(\overline{\tau})_{\gamma}$, is the subset of $M_{I\times J}(R)$ consisting of the matrices whose $(i,j)$-entry belongs to $R_{\sigma_{i}\gamma\tau_{j}^{-1}}$ for all $i\in I,j\in J$. Note that $\sum_{\gamma\in\Gamma}\operatorname{M}_{I\times J}(R)(\overline{\sigma})(\overline{\tau})_{\gamma}$ is, in fact, a direct sum because, for each $i\in I$ and $j\in J$, the sum $\sum_{\gamma\in\Gamma}R_{\sigma_{i}\gamma\tau_{j}^{-1}}$ is direct. Furthermore, if $\gamma\in\Gamma$ and $0\neq a\in R_{\gamma}$, then, for each $i\in I$ and $j\in J$, $aE_{ij}\in\operatorname{M}_{I\times J}(R)(\overline{\sigma})(\overline{\tau})_{\sigma_{i}^{-1}\gamma\tau_{j}}$. Therefore, this indeed defines a grading of $\operatorname{M}_{I\times J}(R)$ by $\Gamma$. Note that
@@
If $I$ and $J$ are finite with $|I|=m,|J|=n\in\mathbb{Z}_{>0}$, we will also write $\operatorname{M}_{m\times n}(R)$, and $\operatorname{M}_{m\times n}(R)(\overline{\sigma})(\overline{\tau})$. If $K$ is a nonempty subset of $I$, we will write $\overline{\sigma}_{K}$ to denote the sequence $(\sigma_{i})_{i\in K}\in(e\Gamma)^{K}$.
@@
Let $I$ be a nonempty set, $e_{0}\in\Gamma_{0}$ such that $\operatorname{supp}(R)\subseteq e_{0}\Gamma e_{0}$ and $\overline{\sigma}=(\sigma_{i})_{i\in I}\in(e_{0}\Gamma)^{I}$ be a $d$-finite sequence. For each $e\in\Gamma_{0}$, consider the finite set $I_{e}:=\{i\in I:d(\sigma_{i})=e\}$. Then, for all $e,f,g\in\Gamma_{0}$, $\gamma\in e\Gamma f$ and $\delta\in f\Gamma g$, there exist isomorphisms of additive groups
@@
[…]
@@
that associates to a matrix of degree $\gamma$, the matrix obtained considering just the entries in $I_{e}\times I_{f}$ because all other entries are zero. Then $\phi_{e,f}:=\bigoplus_{\gamma\in e\Gamma f}\phi_{\gamma}$ is well defined and a gr-isomorphism of $\Gamma$-graded additive groups. The commutativity of the diagram follows by the way matrices are multiplied. ∎
@@ 4.1. General facts on graded division rings
We say that $D$ is a gr-domain or a graded domain if $D\neq\{0\}$ and for all $\gamma,\delta\in\Gamma$ with $d(\gamma)=r(\delta)$ and nonzero elements $a\in D_{\gamma}$, $b\in D_{\delta}$, we have $ab\neq 0$. We say that $D$ is a gr-division ring or a graded division ring if $D\neq\{0\}$ and for all $\gamma\in\operatorname{supp}D$ and nonzero $a\in D_{\gamma}$, there exists an element $a^{-1}\in D_{\gamma^{-1}}$ such that $aa^{-1}=1_{r(\gamma)}$ and $a^{-1}a=1_{d(\gamma)}$. In [34, Section 2], the element $a^{-1}$ is called the $\Gamma$-inverse of $a$, but we prefer to say that $a^{-1}$ is the inverse of $a$. We will also write that $a$ is invertible (with inverse $a$). Note that no confusion will arise with the usual concept of invertibility in ring theory because we will not deal with it. A gr-division ring is a gr-domain. Indeed, given $\gamma,\delta\in\Gamma$ with $d(\gamma)=r(\delta)$ and elements $a\in D_{\gamma}$, $b\in D_{\delta}$, suppose that $ab=0$. If $b\neq 0$, then $0=abb^{-1}=a1_{d(\gamma)}=a$.
@@
Let $D$ be a $\Gamma$-graded division ring. The following assertions hold true
@@
    If we define
    $\Gamma_{[e]}=\left\{\gamma\in\Gamma\mid d(\gamma),r(\gamma)\in[e]\right\},\quad D_{[e]}=\bigoplus\limits_{\gamma\in\Gamma_{[e]}}D_{\gamma}$
    for each equivalence class $[e]\in\Gamma_{0}^{\prime}/\sim$, then $D_{[e]}$ is a nonzero graded ideal of $D$ and $D_{[e]}D_{[f]}=\{0\}$ for $[e]\neq[f]\in\Gamma_{0}^{\prime}/\sim$.
@@
    $D=\bigoplus\limits_{[e]\in\Gamma_{0}/\sim}D_{[e]}$ and $D_{[e]}$ is a gr-simple gr-division ring for each equivalence class $[e]\in\Gamma_{0}^{\prime}/\sim$
@@
(2) It is enough to show that $D_{[e]}D_{[e]}\subseteq D_{[e]}$ and $D_{[e]}D_{[f]}=0$ for all $e,f\in\Gamma^{\prime}_{0}$ with $[e]\neq[f]$. Thus, fix different $[e],[f]\in{\Gamma^{\prime}_{0}}/{\sim}$. Given $\gamma,\gamma^{\prime}\in\Gamma$ such that $r(\gamma),d(\gamma),r(\gamma^{\prime}),d(\gamma^{\prime})\in[e]$, we have $D_{\gamma}D_{\gamma^{\prime}}\subseteq D_{\gamma\gamma^{\prime}}\subseteq D_{[e]}$. This shows $D_{[e]}D_{[e]}\subseteq D_{[e]}$. Now let $\delta\in\Gamma$ be such that $r(\delta),d(\delta)\in[f]$. If we had $D_{\gamma}D_{\delta}\neq 0$, we would have $D_{\gamma\delta}\neq 0$ and therefore, $1_{r(\gamma)}D1_{d(\delta)}\neq 0$ and consequently $e\sim r(\gamma)\sim d(\delta)\sim f$, a contradiction. Hence, $D_{\gamma}D_{\delta}=0$ for all $\gamma,\delta\in\Gamma$ such that $r(\gamma),d(\gamma)\in[e]$ and $r(\delta),d(\delta)\in[f]$. That is, $D_{[e]}D_{[f]}=0$.
@@
(4) Follows from (2) and (3). ∎
There exists a way of describing gr-prime components of a gr-division ring as crossed products. It can be done as in [34, Proposition 2.7] with maps
@@
=$=\left\{\begin{array}[]{ll}0&\textrm{if }d(\gamma)\neq r(\delta)\\ h_{a}h_{b}E_{r(\gamma)d(\delta)}&\textrm{if }d(\gamma)=r(\delta)\end{array}\right.$
=()$\stackrel{{\scriptstyle(*)}}{{=}}\left\{\begin{array}[]{ll}0&\textrm{if }d(\gamma)\neq r(\delta)\\ h_{ab}E_{r(\gamma\delta)d(\gamma\delta)}&\textrm{if }d(\gamma)=r(\delta)\end{array}\right.$
@@
Moreover, for each $f\in\Gamma_{0}^{\prime}(D)$, $\Phi(1_{f})=u_{f}1_{f}u_{f}^{-1}E_{ff}=1_{e}E_{ff}=\mathbb{I}_{f}$. Therefore $\Phi$ is a gr-homomorphism of rings. It remains to show that $\Phi$ is bijective. For that, we construct an inverse $\Psi\colon\operatorname{M}_{\Gamma^{\prime}_{0}(D)}(H)(\overline{\sigma})\rightarrow D$ as follows. Note that if $\gamma\in\operatorname{supp}\operatorname{M}_{\Gamma^{\prime}_{0}(D)}(H)(\overline{\sigma})$, then $\sigma_{f}\gamma\sigma_{f^{\prime}}^{-1}$ is defined in $\Gamma$ for some $f,f^{\prime}\in\Gamma^{\prime}_{0}(D)$. But this happens if, and only if, $f=r(\gamma)$ and $f^{\prime}=d(\gamma)$, i.e., $r(\gamma),d(\gamma)\in\Gamma^{\prime}_{0}(D)$. If $h\in D_{\sigma_{r(\gamma)}\gamma\sigma_{d(\gamma)}^{-1}}$, then $hE_{r(\gamma)d(\gamma)}\in\operatorname{M}_{\Gamma^{\prime}_{0}(D)}(H)(\overline{\sigma})_{\gamma}$ and we define $\Psi(hE_{r(\gamma)d(\gamma)})=u_{r(\gamma)}^{-1}hu_{d(\gamma)}\in D_{\gamma}$.

### v2

@@ 3.6. Some isomorphisms between categories of graded modules
is an isomorphism of groupoids, see for example [9, p.125]. Of course, if $\Gamma$ is a groupoid of the form $I\times G\times I$ for some group $G$ and set $I$, there is a natural choice of isomorphism in (3.1) making $\sigma_{e}=(e_{0},1_{G},e)$ for all $e\in I\equiv\Gamma_{0}$. We remark on passing that we obtain a relation between rings graded by a connected groupoid and categories graded by a group from Example 2.3(2) and (3.1).
Keep in mind the context of the previous paragraph and let $R=\bigoplus_{\gamma\in\Gamma}R_{\gamma}$ be a $\Gamma$-graded ring. We proceed to define the category of right $R$-modules graded by the group $G$, as well as the category of right $R$-modules graded by $G\times\Gamma_{0}$ (both notions will be explicitly defined below). We then relate these categories with $e\Gamma-\operatorname{gr\textrm{-}\!}R$ in Proposition 3.20. Later, in Corollary 5.42, we will use this Proposition to relate the gr-semisimplicity of the $\Gamma$-graded ring $R$ with that of the objects in those categories.
The objects of the category $G-\operatorname{gr\textrm{-}\!}R$ are the $G$-graded $R$-modules. That is, right $R$-modules $M$ such that $MR=M$ and for which there exists a family $\{M_{g}\colon g\in G\}$ of additive subgroups of $M$ such that $M=\bigoplus_{g\in G}M_{g}$ as additive groups and $M_{g}R_{\gamma}\subseteq M_{gg_{\gamma}}$ for each $\gamma\in\Gamma$. If $N$ is another $G$-graded right $R$-module, a homomorphism of $G$-graded modules is a homomorphism of modules $p\colon M\rightarrow N$ such that $p(M_{g})\subseteq N_{g}$ for all $g\in G$. An example of a $G$-graded $R$-module is $R(e)$, where $e\in\Gamma_{0}$, via $R(e)_{g}:=\bigoplus_{f\in\Gamma_{0}}R_{\sigma_{e}^{-1}g\sigma_{f}}$ for each $g\in G$.
The objects of the category $(G\times\Gamma_{0})-\operatorname{gr\textrm{-}\!}R$ are the $(G\times\Gamma_{0})$-graded $R$-modules $M$. That is, right $R$-modules $M$ such that $MR=M$ and for which there exists a family of additive subgrups $\{M_{(g,e)}\colon(g,e)\in G\times\Gamma_{0}\}$ such that $M=\bigoplus\limits_{(g,e)\in G\times\Gamma_{0}}M_{(g,e)}$ as additive groups and, for each $\gamma\in\Gamma$, $M_{(g,e)}R_{\gamma}\subseteq M_{(gg_{\gamma},d(\gamma))}$ if $r(\gamma)=e$ and $M_{(g,e)}R_{\gamma}=0$ otherwise. Given $(G\times\Gamma_{0})$-graded $R$-modules $M,N$, a homomorphism of $(G\times\Gamma_{0})$-graded modules is a homomorphism of $R$-modules $p\colon M\rightarrow N$ such that $p(M_{(g,e)})\subseteq N_{(g,e)}$ for all $(g,e)\in G\times\Gamma_{0}$. An example of a $(G\times\Gamma_{0})$-graded $R$-module is $R(e)$, where $e\in\Gamma_{0}$, via $R(e)_{(g,f)}:=R_{\sigma_{e}^{-1}g\sigma_{f}}$ for each $(g,f)\in G\times\Gamma_{0}$.
@@
A $G$-graded right $R$-module $M=\bigoplus_{g\in G}M_{g}$ has a natural structure of $(G\times\Gamma_{0})$-graded module. Indeed, for each $g\in G$ and $f\in\Gamma_{0}$, consider the identity element $1_{f}\in R_{f}$ and define
@@ 3.7. Gradings on retangular matrix groups
Let $I,J$ be non-empty sets. We denote by $\operatorname{M}_{I\times J}(R)$ the additive group consisting of the $I\times J$ matrices with entries in $R$ and with at most a finite number of nonzero entries. Suppose, moreover, that $\operatorname{supp}(R)\subseteq e\Gamma e$ for some $e\in\Gamma_{0}$. If $\overline{\sigma}=(\sigma_{i})_{i\in I}\in(e\Gamma)^{I},$ $\overline{\tau}=(\tau_{j})_{j\in J}\in(e\Gamma)^{J}$. We will denote by $\operatorname{M}_{I\times J}(R)(\overline{\sigma})(\overline{\tau})$ the $\Gamma$-graded additive group whose homogeneous component of degree $\gamma\in\Gamma$, $\operatorname{M}_{I\times J}(R)(\overline{\sigma})(\overline{\tau})_{\gamma}$, is the subset of $M_{I\times J}(R)$ consisting of the matrices whose $(i,j)$-entry belongs to $R_{\sigma_{i}\gamma\tau_{j}^{-1}}$ for all $i\in I,j\in J$. Note that $\sum_{\gamma\in\Gamma}\operatorname{M}_{I\times J}(R)(\overline{\sigma})(\overline{\tau})_{\gamma}$ is, in fact, a direct sum because, for each $i\in I$ and $j\in J$, the sum $\sum_{\gamma\in\Gamma}R_{\sigma_{i}\gamma\tau_{j}^{-1}}$ is direct. Furthermore, if $\gamma\in\Gamma$ and $0\neq a\in R_{\gamma}$, then, for each $i\in I$ and $j\in J$, $aE_{ij}\in\operatorname{M}_{I\times J}(R)(\overline{\sigma})(\overline{\tau})_{\sigma_{i}^{-1}\gamma\tau_{j}}$. Therefore, this indeed defines a grading of $\operatorname{M}_{I\times J}(R)$ by $\Gamma$. Note that
@@
If $I$ and $J$ are finite with $|I|=m,|J|=n\in\mathbb{Z}_{>0}$, we will also write $\operatorname{M}_{m\times n}(R)$, and $\operatorname{M}_{m\times n}(R)(\overline{\sigma})(\overline{\tau})$. If $K$ is a non-empty subset of $I$, we will write $\overline{\sigma}_{K}$ to denote the sequence $(\sigma_{i})_{i\in K}\in(e\Gamma)^{K}$.
@@
Let $I$ be a non-empty set, $e_{0}\in\Gamma_{0}$ such that $\operatorname{supp}(R)\subseteq e_{0}\Gamma e_{0}$ and $\overline{\sigma}=(\sigma_{i})_{i\in I}\in(e_{0}\Gamma)^{I}$ be a $d$-finite sequence. For each $e\in\Gamma_{0}$, consider the finite set $I_{e}:=\{i\in I:d(\sigma_{i})=e\}$. Then, for all $e,f,g\in\Gamma_{0}$, $\gamma\in e\Gamma f$ and $\delta\in f\Gamma g$, there exist isomorphisms of additive groups
@@
${\operatorname{M}_{I}(R)(\overline{\sigma})_{\gamma}\times\operatorname{M}_{I}(R)(\overline{\sigma})_{\delta}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\phi_{\gamma},\phi_{\delta})}$${\operatorname{M}_{|I_{e}|\times|I_{f}|}(R)(\overline{\sigma}_{I_{e}})(\overline{\sigma}_{I_{f}})_{\gamma}\times\operatorname{M}_{|I_{f}|\times|I_{g}|}(R)(\overline{\sigma}_{I_{f}})(\overline{\sigma}_{I_{g}})_{\delta}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{M}_{I}(R)(\overline{\sigma})_{\gamma\delta}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{\gamma\delta}}$${\operatorname{M}_{|I_{e}|\times|I_{g}|}(R)(\overline{\sigma}_{I_{e}})(\overline{\sigma}_{I_{g}})_{\gamma\delta}}$
@@
that associates to a matrix of degree $\gamma$, the matrix obtained considering just the entries in $I_{e}\times I_{f}$ because all other entries are zero. Then $\phi_{e,f}:=\bigoplus_{\gamma\in e\Gamma f}\phi_{\gamma}$ is well-defined and a gr-isomorphism of $\Gamma$-graded additive groups. The commutativity of the diagram follows by the way matrices are multiplied. ∎
@@ 4.1. General facts on graded division rings
We say that $D$ is a gr-domain or a graded domain if $D\neq\{0\}$ and for all $\gamma,\delta\in\Gamma$ with $d(\gamma)=r(\delta)$ and nonzero elements $a\in D_{\gamma}$, $b\in D_{\delta}$, we have $ab\neq 0$. We say that $D$ is a gr-division ring or a graded division ring if $D\neq\{0\}$ and for all $\gamma\in\operatorname{supp}D$ and nonzero $a\in D_{\gamma}$, there exists an element $a^{-1}\in D_{\gamma^{-1}}$ such that $aa^{-1}=1_{r(\gamma)}$ and $a^{-1}a=1_{d(\gamma)}$. In [37, Section 2], the element $a^{-1}$ is called the $\Gamma$-inverse of $a$, but we prefer to say that $a^{-1}$ is the inverse of $a$. We will also write that $a$ is invertible (with inverse $a$). Note that no confusion will arise with the usual concept of invertibility in ring theory because we will not deal with it. A gr-division ring is a gr-domain. Indeed, given $\gamma,\delta\in\Gamma$ with $d(\gamma)=r(\delta)$ and elements $a\in D_{\gamma}$, $b\in D_{\delta}$, suppose that $ab=0$. If $b\neq 0$, then $0=abb^{-1}=a1_{d(\gamma)}=a$.
@@
Let $D$ be a $\Gamma$-graded division ring. The following assertions hold:
@@
    If we define Γ_[e]={γΓd(γ),r(γ)[e]}, D_[e]=Γ_[e] D_γ for each equivalence class $e]\in\Gamma_{0}^{\prime}/\sim$, then $D_{[e]}$ is a nonzero graded ideal of $D$ and $D_{[e]}D_{[f]}=\{0\}$ for $e]\neq[f]\in\Gamma_{0}^{\prime}/\sim$.
@@
    $D=\bigoplus\limits_{[e]\in\Gamma_{0}/\sim}D_{[e]}$ and $D_{[e]}$ is a gr-simple gr-division ring for each equivalence class $e]\in\Gamma_{0}^{\prime}/\sim$
@@
(2) It is enough to show that $D_{[e]}D_{[e]}\subseteq D_{[e]}$ and $D_{[e]}D_{[f]}=0$ for all $e,f\in\Gamma^{\prime}_{0}$ with $e]\neq[f]$. Thus, fix different $e],[f]\in{\Gamma^{\prime}_{0}}/{\sim}$. Given $\gamma,\gamma^{\prime}\in\Gamma$ such that $r(\gamma),d(\gamma),r(\gamma^{\prime}),d(\gamma^{\prime})\in[e]$, we have $D_{\gamma}D_{\gamma^{\prime}}\subseteq D_{\gamma\gamma^{\prime}}\subseteq D_{[e]}$. This shows $D_{[e]}D_{[e]}\subseteq D_{[e]}$. Now let $\delta\in\Gamma$ be such that $r(\delta),d(\delta)\in[f]$. If we had $D_{\gamma}D_{\delta}\neq 0$, we would have $D_{\gamma\delta}\neq 0$ and therefore, $1_{r(\gamma)}D1_{d(\delta)}\neq 0$ and consequently $e\sim r(\gamma)\sim d(\delta)\sim f$, a contradiction. Hence, $D_{\gamma}D_{\delta}=0$ for all $\gamma,\delta\in\Gamma$ such that $r(\gamma),d(\gamma)\in[e]$ and $r(\delta),d(\delta)\in[f]$. That is, $D_{[e]}D_{[f]}=0$.
@@
[…]
There exists a way of describing gr-prime components of a gr-division ring as crossed products. It can be done as in [37, Proposition 2.7] with maps
@@
$=\left\{\begin{array}[]{ll}0&\textrm{if }d(\gamma)\neq r(\delta)\\ h_{a}h_{b}E_{r(\gamma)d(\delta)}&\textrm{if }d(\gamma)=r(\delta)\end{array}\right$.
$\mathrel{{\mathop{=}\limits^{(*)}}}\left\{\begin{array}[]{ll}0&\textrm{if }d(\gamma)\neq r(\delta)\\ h_{ab}E_{r(\gamma\delta)d(\gamma\delta)}&\textrm{if }d(\gamma)=r(\delta)\end{array}\right$.
@@
Moreover, for each $f\in\Gamma_{0}^{\prime}(D)$, $\Phi(1_{f})=u_{f}1_{f}u_{f}^{-1}E_{ff}=1_{e}E_{ff}=\mathbb{I}_{f}$, the identity element of $\operatorname{M}_{\Gamma_{0}^{\prime}(D)}(H)(\overline{\sigma})_{f}$. Therefore $\Phi$ is a gr-homomorphism of rings. It remains to show that $\Phi$ is bijective. For that, we construct an inverse $\Psi\colon\operatorname{M}_{\Gamma^{\prime}_{0}(D)}(H)(\overline{\sigma})\rightarrow D$ as follows. Note that if $\gamma\in\operatorname{supp}\operatorname{M}_{\Gamma^{\prime}_{0}(D)}(H)(\overline{\sigma})$, then $\sigma_{f}\gamma\sigma_{f^{\prime}}^{-1}$ is defined in $\Gamma$ for some $f,f^{\prime}\in\Gamma^{\prime}_{0}(D)$. But this happens if and only if $f=r(\gamma)$ and $f^{\prime}=d(\gamma)$, i.e., $r(\gamma),d(\gamma)\in\Gamma^{\prime}_{0}(D)$. If $h\in D_{\sigma_{r(\gamma)}\gamma\sigma_{d(\gamma)}^{-1}}$, then $hE_{r(\gamma)d(\gamma)}\in\operatorname{M}_{\Gamma^{\prime}_{0}(D)}(H)(\overline{\sigma})_{\gamma}$ and we define $\Psi(hE_{r(\gamma)d(\gamma)})=u_{r(\gamma)}^{-1}hu_{d(\gamma)}\in D_{\gamma}$.